I need help solving this problem:
The grocery store we use does not mark prices on its goods. my wife went to this store and bought 5 packages of bacon and 3 cartons of eggs and paid a total of 18.52. i also went to the store and bought 3 packages of bacon and 5 cartons of eggs and paid a total of 14.92. now we want to return 3 packages of bacon and 3 cartons of eggs. how much will be refunded?

Answer by mananth(16946)

The grocery store we use does not mark prices on its goods. my wife went to this store and bought 5 packages of bacon and 3 cartons of eggs and paid a total of 18.52. i also went to the store and bought 3 packages of bacon and 5 cartons of eggs and paid a total of 14.92. now we want to return 3 packages of bacon and 3 cartons of eggs. how much will be refunded?
Let a packet of bacon cost $x
Let a packet of eggs cost $ y
wife: 5x+3y= 18.52
Yourself
3x+5y= 14.92
3(5x+3y)-5(3x+5y)= 3*18.52-5*14.92
15x+9y-15x-25y= 55.56-74.6
-16y=19.04
x=1.19 cost of 1 packet of bacon
3x+5y=14.92
3*1.19+5y=14.92
5y= 14.92-3.57
y= 2.27 cost of 1 packet of eggs
3 packets of bacon = 3*1.19= 3.57
3 packages of eggs = 3*2.27= 6.81
you should get a refund of $ 10.38